OpenStudy (anonymous):

Air is pumped into a spherical balloon at a rate of 8cm^3 per second. At what rate is the radius of the sphere changing when its volume is 36 pie cm^3?

OpenStudy (anonymous):

related rate problem right? hold on a second i am in the middle of another problem, but the idea is to write the equation for the volume of the sphere in terms of the radius, differentiate wrt r and then substitute 8 for \(V'\) and 36 for \(V\)

OpenStudy (anonymous):

you got the formula for the volume of the sphere?

OpenStudy (anonymous):

yes it is v=(4/3)[pi\]r^3

OpenStudy (anonymous):

\[V=\frac{2}{3}\pi r^3\] take the derivative of both sides wrt \(t\) and get \[V'=2\pi r^2 r'\] now make the replacements and solve for \(r'\)

OpenStudy (anonymous):

ok that was wrong, it is \[V=\frac{4}{3}\pi r^3\]so \[V'=4\pi r^2 r'\]

OpenStudy (anonymous):

so i will plug in 8 and get 8=4[pi \]r^2r'

OpenStudy (anonymous):

then i would solve for r' but how would i find the radius?

OpenStudy (anonymous):

you know that \[V=\frac{4}{3}\pi r^3\]and you are asked for \(r'\) when \(V=36\) so solve \(36=\frac{4}{3}\pi r^3\) for \(r\)

OpenStudy (anonymous):

okay thank you. the answer would be 2/9pi